OpenStudy (anonymous):

y varies directly as the square of x. When x = 2, y = 40. Find x when y = 640. x = ±6 x = ±8 x = ±12 x = ±20

OpenStudy (shubhamsrg):

it means y = k x^2 (for some constant k) => y1/y2 = (x1/x2)^2 here y1= 40,x1 =2 and y2 = 640, x2= ?

OpenStudy (anonymous):

12?

OpenStudy (shubhamsrg):

try again..

OpenStudy (anonymous):

8

OpenStudy (shubhamsrg):

yep..the very same ^_^ note its both 8 and -8

OpenStudy (anonymous):

Ok, Thanks!